OpenStudy (anonymous):

Points Y and Z are the same distance from point X on a number line. The coordinate of point X is -7 and the coordinate of point Y is -1. What is the coordinate of point Z?

OpenStudy (anonymous):

@blurbendy << smartie

OpenStudy (blurbendy):

if y is -1 and x is -7, then y is 6 units from 7. that means Z is also 6 units from 7. so, what is Z?

OpenStudy (anonymous):

-1

OpenStudy (blurbendy):

no, Y is already -1, go the other way

OpenStudy (anonymous):

1 or ( idk)

OpenStudy (blurbendy):

-7 + 6 = -1 = Y -7 - 6 = ????

OpenStudy (anonymous):

-13

OpenStudy (blurbendy):

yes, good job

OpenStudy (anonymous):

yea ik im slow

OpenStudy (anonymous):

t noon, the atmospheric pressure at a weather station was 763 mmHg. At 4 p.m., the atmospheric pressure had changed by 29 mmHg. Which expressions could represent the atmospheric pressure at the weather station at 4 p.m.? Choose two answers that are correct. A. (29 – 763) mmHg B. –(763 + 29) mmHg C. (763 – 29) mmHg D. (763 + 29) mmHg

OpenStudy (blurbendy):

the atmospheric pressure either increased or decreased. which answers tell you that?

OpenStudy (anonymous):

c or b??

OpenStudy (blurbendy):

c is correct, but b has a - sign in front. D is the other one

OpenStudy (anonymous):

thanks

OpenStudy (anonymous):

A pizzeria is 11 blocks due east from the town center. Emily is 4 blocks away from the pizzeria. Which statement could correctly describe Emily’s location? A. Emily is 4 blocks due east from the town center. B. Emily is 7 blocks due east from the town center. C. Emily is 7 blocks due west from the town center. D. Emily is 11 blocks due west from the town center.

OpenStudy (blurbendy):

think of a numbr line 11 is where the pizzeria is, emily is on 7 or 15 because she is 4 blocks away so, the town center must be on 0. so, which statement is correct?

OpenStudy (anonymous):

its c or b??

OpenStudy (blurbendy):

b because everything in this question is east, not west

OpenStudy (anonymous):

Points R, S, and T are plotted on a number line. The coordinate of point R is –6. Points S and T are the same distance from point R. What could be the coordinates of S and T? A. S = 0, T = 6 B. S = –4, T = –2 C. S = –9, T = –12 D. S = –10, T = –2

OpenStudy (blurbendy):

which answer has units that are the same distance away from -6 ?

OpenStudy (anonymous):

a

OpenStudy (blurbendy):

nope, 6 is 12 units away and 0 is 6 units away

OpenStudy (anonymous):

D

OpenStudy (blurbendy):

yes, good job

OpenStudy (anonymous):

you should teach math

OpenStudy (blurbendy):

lol, i like 2 help!

OpenStudy (anonymous):

you got 100

OpenStudy (blurbendy):

you did, I only helped guide you

OpenStudy (anonymous):

i have 10 more Questions lol ima start a new thing